# Hard View Selection for Self-Supervised Learning

### v4

@@ Appendix H Computational Overhead of HVS
  • switching between HVS and the standard pipepline in alternating fashion or
  • bypassing forwarding of similar pairs and more
Appendix I Hard View Selection Objectives
I.1 SimCLR
In this section, we are going to introduce the application of HVS with the SimCLR objective. Assume a given set of images $\mathcal{D}$, an image augmentation distribution $\mathcal{T}$, a minibatch of $M$ images $\mathbf{x}=\left\{x_{i}\right\}_{i=1}^{M}$ sampled uniformly from $\mathcal{D}$, and two sets of randomly sampled image augmentations $A=\left\{t_{i}\sim\mathcal{T}\right\}_{i=1}^{M}$ and $B$ sampled from $\mathcal{T}$. We apply $A$ and $B$ to each image in $\mathbf{x}$ resulting in $\mathbf{x}^{A}$ and $\mathbf{x}^{B}$. Both augmented sets of views are subsequently projected into an embedding space with $\mathbf{z}^{A}=g_{\theta}(f_{\theta}(\mathbf{x}^{A}))$ and $\mathbf{z}^{B}=g_{\theta}(f_{\theta}(\mathbf{x}^{B}))$ where $f_{\theta}$ represents an encoder (or backbone) and $g_{\theta}$ a projector network. Contrastive learning algorithms then minimize the following objective function:
@@ I.1 SimCLR
To further enhance the training process, we introduce a modification to the loss function where instead of having two sets of augmentations $A$ and $B$, we now have ”N” sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$. Each set $A_{i}$ is sampled from the image augmentation distribution $\mathcal{T}$, and applied to each image in $\mathbf{x}$, resulting in ”N” augmented sets of views $\mathbf{x}^{A_{1}},\mathbf{x}^{A_{2}},\ldots,\mathbf{x}^{A_{N}}$.
@@
Appendix J Attention Maps
In this section, we visualize attention maps as a way to qualitatively study the potential effects and differences of the HVS learning method and their resulting features. The provided attention maps are from a DINO ViT-S/16 100 model, and we contrast the attention for each HVS and the baseline. All input images are from the ImageNet-1k validation split. The color code used depicts strong attention in yellow and weak attention in green and attentions from the HVS model are shown in the top row and attentions from the baseline in the bottom row, respectively. To summrize this study, we do not see apparent strong differences between HVS and the baseline. What we occasionally notice is that the HVS models seem to capture the shape of some subtle/indistinct objects better (e.g. the creek/river in the valley in Fig. 12, the lizard in Fig. 13, or the speaker in Fig. 14 or focus slightly more on the context and background (Fig. 16, 17, and 18).
Figure 11:
Figure 12:
Figure 13:
Figure 14:
Figure 15:
Figure 16:
Figure 17:
Figure 18:
Figure 19:

### v6

@@ I.3 Object Detection and Instance Segmentation
  • other approaches to derive manual augmentation policies or
  • bypassing forwarding of similar pairs.
Appendix K Hard View Pretraining Objectives
K.1 SimCLR
In this section, we are going to introduce the application of HVP with the SimCLR objective. Assume a given set of images $\mathcal{D}$, an image augmentation distribution $\mathcal{T}$, a minibatch of $M$ images $\mathbf{x}=\left\{x_{i}\right\}_{i=1}^{M}$ sampled uniformly from $\mathcal{D}$, and two sets of randomly sampled image augmentations $A=\left\{t_{i}\sim\mathcal{T}\right\}_{i=1}^{M}$ and $B$ sampled from $\mathcal{T}$. We apply $A$ and $B$ to each image in $\mathbf{x}$ resulting in $\mathbf{x}^{A}$ and $\mathbf{x}^{B}$. Both augmented sets of views are subsequently projected into an embedding space with $\mathbf{z}^{A}=g_{\theta}(f_{\theta}(\mathbf{x}^{A}))$ and $\mathbf{z}^{B}=g_{\theta}(f_{\theta}(\mathbf{x}^{B}))$ where $f_{\theta}$ represents an encoder (or backbone) and $g_{\theta}$ a projector network. Contrastive learning algorithms then minimize the following objective function:
@@
To further enhance the training process, we introduce a modification to the loss function where instead of having two sets of augmentations $A$ and $B$, we now have "N" sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$. Each set $A_{i}$ is sampled from the image augmentation distribution $\mathcal{T}$, and applied to each image in $\mathbf{x}$, resulting in "N" augmented sets of views $\mathbf{x}^{A_{1}},\mathbf{x}^{A_{2}},\ldots,\mathbf{x}^{A_{N}}$.